OpenStudy (anonymous):

how do you solve: 6x+2=4x+14 ?

OpenStudy (anonymous):

6x - 4x = 14 -2 2x = 12 x = 12/2 x = 6

OpenStudy (anonymous):

With Algebra! Hoorah! Diogo already solved it for you but in the future, for questions like this, think first about moving all the terms with an "x" or variable in them to one side. In other words, isolate the variable. You can see Diogo did that as his first step.

OpenStudy (anonymous):

Thanks!